### v1

# Collaborative Learning in Agentic Systems: A Collective AI is Greater Than the Sum of Its Parts

Abstract
Agentic AI has gained significant interest as a research paradigm focused on autonomy, self-directed learning, and long-term reliability of decision making. Real-world agentic systems operate in decentralized settings on a large set of tasks or data distributions with constraints such as limited bandwidth, asynchronous execution, and the absence of a centralized model or even common objectives. We posit that exploiting previously learned skills, task similarities, and communication capabilities in a collective of agentic AI are challenging but essential elements to enabling scalability, open-endedness, and beneficial collaborative learning dynamics. In this paper, we introduce Modular Sharing and Composition in Collective Learning (MOSAIC), an agentic algorithm that allows multiple agents to independently solve different tasks while also identifying, sharing, and reusing useful machine-learned knowledge, without coordination, synchronization, or centralized control. MOSAIC combines three mechanisms: (1) modular policy composition via neural network masks, (2) cosine similarity estimation using Wasserstein embeddings for knowledge selection, and (3) asynchronous communication and policy integration. Results on a set of RL benchmarks show that MOSAIC has a greater sample efficiency than isolated learners, i.e., it learns significantly faster, and in some cases, finds solutions to tasks that cannot be solved by isolated learners. The collaborative learning and sharing dynamics are also observed to result in the emergence of ideal curricula of tasks, from easy to hard. These findings support the case for collaborative learning in agentic systems to achieve better and continuously evolving performance both at the individual and collective levels.
1 Introduction
The ability to handle multiple dynamic tasks concurrently and over long-term time windows De Lange et al. (2022); Parisi et al. (2019) is a coveted feature of autonomous, adaptive agents that operate in open-ended environments involving physical-, vision-, and language-based tasks. Dynamic environments demand resilient agents that continuously learn useful behaviors from local data, generalize across diverse tasks, and effectively share useful knowledge (Tarale et al., 2025). Such requirements are central to the emerging vision of agentic AI systems, which emphasize self-directed learning, adaptation, asynchronous operation, and scalable autonomy (Colas et al., 2022; Forestier et al., 2022). Recent work on collaborative and collective AI further highlights the importance of decentralized learning, peer-to-peer sharing, and modular reuse for building scalable societies of AI agents (Soltoggio et al., 2024).
In such scenarios, reinforcement learning (RL) signals that can be sourced from the environment allow for policy optimization without external supervision, making RL an effective tool used both in robotics and language models (Sutton and Barto, 2018; Jaderberg et al., 2017; Wang et al., 2025; Sun, 2023). Crucially, knowledge reuse has been a central assumption in multi-task (MTRL) and meta reinforcement learning (Meta-RL) (Hospedales et al., 2021), in which learning of many tasks is facilitated by exploiting similarities across tasks. The successful deployment of pre-trained models in LLMs is another example of effective reuse of previously acquired knowledge across tasks Han et al. (2021); Qiu et al. (2020). Reliability of RL algorithms when dealing with evolving streams of data is improved with lifelong RL approaches (Ben-Iwhiwhu et al., 2023).
Most existing MTRL, Meta-RL, and lifelong RL methods assume a centralized setup with shared policy networks and known task labels (Wang et al., 2023; Sun et al., 2022; Hendawy et al., 2024; Amani et al., 2023; Zeng et al., 2021; D’Eramo et al., 2020). These assumptions limit their applicability in fully decentralized or real-world agentic scenarios, where agents typically operate independently, and communication, when possible, is asynchronous and not coordinated. To address these realistic conditions, recent work in lifelong learning and multi-agent systems explored machine-learned knowledge sharing among many agents (Nath et al., 2023; Rostami et al., 2017; Gao et al., 2021; Ge et al., 2023; Gerstgrasser et al., 2023). Similar objectives are pursued by recent approaches that share and compose LoRA modules in LLMs (Huang et al., 2024; Roy et al., 2025), demonstrating that this challenge is crucial to scale current AI approaches. A full coverage of the related works in the literature can be found in appendix A.
While the idea of sharing and reusing knowledge in agentic machine learning systems is intuitive and appealing, a key algorithmic problem has not yet been well addressed: in a collective of AI agents solving different tasks, what specific knowledge should be shared, with whom, and when? DARPA/MTO (2021); Soltoggio et al. (2024) Moreover, does a decentralized collective of agents have an advantage over isolated agents learning the same tasks?
To address these unanswered questions, this study proposes a new algorithm, Modular Sharing and Composition in Collective Learning (MOSAIC), that describes what policies to exchange among peers, specifically among which agents, when to do so, and how such policies can be integrated. Extending and integrating recent advances in lifelong machine learning, the approach employs modular knowledge representations in the form of network masks and proposes a knowledge exchange mechanism based on cosine similarity computed with Wasserstein task embeddings. Performance measures and similarities are used by individual agents to determine specifically what knowledge to retrieve, from whom, and when.
The experimental evidence and the ablation studies show that precise selection of specific knowledge is essential to exploit the advantages of sharing. This evidence indicates that not all knowledge is beneficial to all agents, and that modular composition and cosine similarity metrics are essential ingredients in a collective of collaborating agents. MOSAIC agents learn faster than isolated agents by selectively leveraging the available knowledge from peers working on similar tasks. In some cases, by combining knowledge retrieved from other agents working on similar tasks, MOSAIC agents are able to solve tasks that are too difficult for isolated learners. Further analysis reveals that the collective implicitly discovers an ideal curriculum in which agents working on simpler tasks subsequently help agents working on harder tasks. The MOSAIC learning dynamics indicate that a collective of collaborative MOSAIC agents can achieve superior performance than that of an equivalent number of isolated learners, making a strong case for targeted knowledge sharing in agentic systems.
1.1 Contributions
This work introduces MOSAIC, an agentic system, i.e., an autonomous decision-making and self-contained individual agent, that implements autonomous and self-directed knowledge exchange with other MOSAIC peers. MOSAIC is a novel integration of the following algorithmic steps: (i) modularized knowledge learning and representation; (ii) online Wasserstein task embeddings computed from reinforcement learning datasets; (iii) knowledge querying and retrieval with cosine similarity (between embeddings) and performance-based selection heuristics.
The performance evaluations reveal significant learning advantages for a collective of collaborating agents with respect to isolated learners, which, in some cases, results in successful policies that cannot be discovered in isolation.
Lastly, the analysis of learning dynamics revealed that agents choose what to learn in ways that lead to the emergence of implicit curricula, task clusters, and task hierarchies. This property enables an open-ended training behavior in which agents autonomously determine their learning trajectories by reusing specific subsets from the available collective knowledge.
2 Methodology
The following section describes the components of the MOSAIC algorithm that includes knowledge representations (Section 2.1), Wasserstein task embedding (Section 2.2), knowledge selection and sharing (Section 2.3), and knowledge composition (Section 2.4). Finally, Section 2.5 illustrates the integration of all components with pseudo-code. A visual summary of the main algorithmic steps is provided in Figure 1.
2.1 Policy Representations
The ability to encapsulate specific knowledge in isolated module representations (Alet et al., 2018) is a key component of MOSAIC. To this aim, a variety of parameter isolation mechanisms from the field of lifelong learning can be used von Oswald et al. (2020); Hu et al. (2022); Rusu et al. (2016); Mallya and Lazebnik (2018). Among those, promising results in lifelong RL were shown by network masks (Ben-Iwhiwhu et al., 2023; Wortsman et al., 2020) that are therefore used in this study. Masks are naturally suited for composition through linear combinations, which is an essential mechanism in MOSAIC. The use of masking makes MOSAIC agnostic to the specific reinforcement learning algorithm used, requiring only that policies are parameterized by a neural network. This allows it to be layered on top of any standard RL method, such as PPO (Schulman et al., 2017), DQN (Mnih et al., 2013), or A3C (Mnih et al., 2016).
Figure 1: High-level illustration of the main MOSAIC algorithmic steps. (A) A Wasserstein task embedding $v_{\tau}$ is computed from a SAR batch (Section 2.2) by Agent A0 (representing any agent in the collective). (B) Periodically, the agent A0 broadcasts a task embedding query (TEQ) to all known peers. (C) Peers send back a query response (QR) that contains their task’s $v_{\tau},r$, and the corresponding mask ID; (D) Agent A0 selects relevant embeddings using cosine similarity on the Wasserstein embeddings. (E) A0 further selects relevant embeddings using the performance criterion (Section 2.3). (F) A0 sends mark requests (MR). (G) The contacted agents respond by sending the requested mask through a mask transfer (MTR). (H) Incoming masks from A2 and A3 are incorporated into A0 (Section 2.4). The training of the agent’s policy occurs in parallel (not represented in the figure).
Preliminaries.
Each task $\tau\in\mathcal{T}$ is modeled as an MDP defined by the tuple $\mathcal{M}=\left\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\right\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{P}(s^{\prime}|s,a)$ is the transition probability function, $\mathcal{R}(s,a)\in\mathbb{R}$ is the reward received after taking action $a$ in the state $s$, and $\gamma\in[0,1]$ is the discount factor. An agent interacts with its environment and aims to learn a policy $\pi_{\tau}:\mathcal{S}\to\mathcal{A}$ that maximizes the expected discounted return,
$J_{\tau}(\pi_{\tau})=\mathbb{E}_{\pi_{\tau}}\left[\sum_{t=0}^{\infty}\gamma^{t}_{\tau}\mathcal{R}_{\tau}(s_{t},a_{t})\right].$ (1)
Mask Representations.
The mask representation used in this study relies on a frozen backbone network $\Phi$ that must be the same for all agents. Each policy $\pi_{\tau}$ is implicitly parameterized through a sparse mask $\phi_{\tau}$, such that $\pi_{\tau}=\pi_{\Phi\odot g(\phi_{\tau})}$, where $\odot$ denotes the Hadamard product, and $g(\cdot)$ is a binarization function applied during the forward pass. The mask $\phi_{\tau}$ consists of a set of real-valued score vectors, one per layer of the shared backbone network $\Phi$, which select a sparse, task-specific subnetwork from $\Phi$ without modifying its parameters. The scores are binarized during the forward pass to produce a hard mask that is used for inference. The binarization is performed during inference, using the element-wise thresholding function,
$g(\phi_{\tau})_{\ell}=\begin{cases}1&\text{if }(\phi_{\tau})_{\ell}>\epsilon\\ 0&\text{otherwise}\end{cases},\quad\text{with }\epsilon=0\quad,$ (2)
where $g(\phi_{\tau})$ denotes the full binary mask and $\ell$ indexes the layers of the backbone network $\Phi$. During the backward pass, gradients are computed with respect to the real-valued mask parameters $\phi_{\tau}$, and updates are performed using the straight-through estimator (STE) (Wortsman et al., 2020) to enable learning through the non-differentiable binarization step.
2.2 Wasserstein Task Embeddings for Online Reinforcement Learning
A key intuition in this study is that knowledge composition is effective if policies from other agents are acquired only when they are likely to be beneficial, and because they have been derived from related tasks. This hypothesis was verified with an ablation study. MOSAIC implements this idea by computing task embeddings from trajectory data and using them to identify semantically related peers via geometric proximity. This idea builds on work in latent task inference and representation learning (Zintgraf et al., 2021; Achille et al., 2019), adopting a non-parametric approach aligned with prototype-based and optimal transport methods (Yarats et al., 2021; Chu et al., 2024; Liu et al., 2022).
Unlike gradient-based adaptation techniques (Lan et al., 2019), which are unstable in RL and costly at inference time, MOSAIC computes task embeddings through training-free distributional comparisons, making them well-suited for decentralized, real-time systems requiring modular policy reuse without supervision. To support these requirements, MOSAIC extends the Wasserstein Task Embedding (WTE) framework of Liu et al. (2022) to the RL domain by defining an online embedding function over batches of state-action-reward (SAR) tuples drawn from an agent’s replay buffer. For task $\tau$, the agent maintains an empirical task distribution $\mu_{\tau}\in\mathcal{P}(\mathbb{R}^{d})$, defined as,
$\mu_{\tau}=\frac{1}{N}\sum_{t=1}^{N}\delta_{x_{t}},\quad\text{where}\quad x_{t}=(s_{t},a_{t},r_{t})\quad\text{and}\quad x_{t}\sim\mathcal{D}_{\tau},$ (3)
where $\mathcal{D}_{\tau}$ denotes the replay buffer for task $\tau$, and each $x_{t}\in\mathbb{R}^{d}$ is a state-action-reward (SAR) tuple. The tuple dimensionality is given by $d=d_{s}+d_{a}+d_{r}$, where $d_{s}$, $d_{a}$, and $d_{r}$ are the dimensions of the state, action, and scalar reward, respectively. In our experiments, we use $N=128$ samples to compute the empirical task distribution. $\delta_{x_{t}}$ denotes the Dirac measure centered at the sample $x_{t}$. A fixed synthetic reference distribution $\mu_{0}$ is defined as,
$\mu_{0}=\frac{1}{M}\sum_{m=1}^{M}\delta_{x_{m}^{0}},$ (4)
where $x_{m}^{0}\sim\text{Uniform}(-1,1)^{d}$ is sampled once during initialization and fixed thereafter. We fix $M=50$ as the number of reference points in the synthetic distribution. This shared reference is used across all agents and enables them to align their embeddings in a shared latent space without centralized training or supervision. The 2-Wasserstein distance between $\mu_{\tau}$ and $\mu_{0}$ is computed by solving the optimal transport problem,
$\gamma^{*}=\arg\min_{\gamma\in\mathbb{R}^{N\times M}}\sum_{t=1}^{N}\sum_{m=1}^{M}\gamma_{tm}\|x_{t}-x^{0}_{m}\|^{2},\quad\text{s.t. }\gamma_{tm}\geq 0,\gamma\mathbf{1}_{M}=\frac{1}{N}\mathbf{1}_{N},\gamma^{\top}\mathbf{1}_{N}=\frac{1}{M}\mathbf{1}_{M},$ (5)
where $\mathbf{1}_{N}$ and $\mathbf{1}_{M}$ are vectors of ones in $\mathbb{R}^{N}$ and $\mathbb{R}^{M}$, respectively. The Wasserstein task embedding operator $\psi$ maps the task distribution $\mu_{\tau}$ to a vector $v_{\tau}\in\mathbb{R}^{M\times d}$ via the barycenter projection of $\gamma^{*}$ on $\{x_{t}\}_{t=1}^{N}$,
$v_{\tau}=\psi(\mu_{\tau})=\left[\sum_{t=1}^{N}\gamma_{tm}^{*}x_{t}\right]_{m=1}^{M}.$ (6)
This mapping places each task $\tau$ in a shared latent space, where metrics can be used to quantify task relationships. A 2-point moving average is used to smooth behavioral fluctuations and stabilize evolving policies in MOSAIC. This embedding strategy enables agents to reason about task-relatedness in constant time, without additional training or forward passes. Importantly, it fits the online and asynchronous nature of agentic RL systems, while remaining robust to evolving policies and lightweight enough for continuous use in edge-deployed systems.
2.3 Knowledge Selection and Sharing
Agents communicate maintaining a list of IP addresses and ports for other agents, which allows them to be located anywhere on the Internet. Each agent $i\in\mathcal{I}$ can directly communicate with all other agents $j\in\mathcal{I},\ j\neq i$ at any time.
Phase 1: Query-Metadata.
At any time, an agent $i$ can initiate a task embedding query (TEQ) by broadcasting its current task embedding $v_{i}$, iteration performance $r_{i}$, and its IP/port (Figure 1(B)). Upon receiving the TEQ, each peer agent $j\neq i$ responds with its most recent task embedding $v_{j}$ and associated performance score $r_{j}$.
Phase 2: Mask Selection and Acquisition.
Once the agent has received query responses (QR) with relevant embeddings from other agents, it computes the cosine similarity between its own embedding and those of each peer,
$d_{\text{cos}}(v_{i},v_{j})=\frac{\langle\text{vec}(v_{i}),\text{vec}(v_{j})\rangle}{|\text{vec}(v_{i})|\cdot|\text{vec}(v_{j})|},\quad d_{\text{cos}}\in[-1,1].$ (7)
In practice, cosine similarities between Wasserstein task embeddings are typically positive. Cosine similarity was observed to perform better than Euclidean distance in preliminary tests, and showed scale-invariance and empirical robustness in high-dimensional settings. To identify useful policies, agent $i$ applies two heuristic filters based on similarity and performance:
$\mathbb{I}_{\text{align}}(i,j)$ =$=\begin{cases}1&\text{if }d_{\text{cos}}(v_{i},v_{j})>\theta,\\ 0&\text{otherwise}\end{cases}$ (Criterion 1)
$\mathbb{I}_{\text{perf}}(i,j)$ =$=\begin{cases}1&\text{if }r_{j}>r_{i},\\ 0&\text{otherwise}\end{cases}$ (Criterion 2)
with $\theta=$ set to $0.5$ in all experiments. The first criterion promotes semantic alignment between tasks, revealing latent similarities that may lead to effective transfer (see Figure 4). The second ensures that agents only acquire masks from peers whose performance exceeds their own, avoiding noisy or under-trained policies that can degrade performance. Criterion 2 guards against the frequent false positives seen when comparing poorly trained policies that yield deceptively similar embeddings. The selection criteria can also be used on the peers after they receive the query. This method acts as a bandwidth-efficient alternative that enables communication without affecting the overall outcome of the approach. Additionally, this method offers a scalable pathway for extensions to lifelong learning, which are discussed in Appendix B.
The peer masks that pass both criteria are stored as the set $P_{c+1}=\{\phi_{1},...,\phi_{K}\}$, where each $\phi_{k}$ denotes a mask received from a selected peer. This two-phase strategy ensures that agents acquire knowledge that is not only relevant but also demonstrably effective.
2.4 Knowledge Composition
Masks received from other agents meeting the two selection criteria above have been pretrained on tasks that are likely related to, or share similarities with, the agent’s current task. As originally shown in Ben-Iwhiwhu et al. (2023), a linear combination of masks, weighted with trainable parameters $\beta$, can lead to beneficial policy search in RL. MOSAIC exploits this idea by combining policies that have been trained on the local agent, plus those received from other agents. In this study, the $\beta$ values are computed via softmax over a set of beta parameters optimized in log-space, $\tilde{\beta}\in\mathbb{R}^{1+|P_{c}|}$, where $P_{c}$ is the number of masks received. During the forward pass, the agent constructs the linearly combined, binarized mask,
$\phi_{\tau}^{\text{lc}}=g\left(\beta_{\tau}\phi_{\tau}+\sum_{k=1}^{|P_{c+1}|}\beta_{k}\phi_{k}\right),$ (8)
where $\phi_{\tau}$ is the agent’s own task mask, and $\phi_{k}$ is the $k$-th peer mask. The parameters $\beta_{\tau}$ and $\beta_{k}$ are the softmax-normalized values derived from $\tilde{\beta}$, for $\phi_{\tau}$ and $\phi_{k}$, respectively. The resulting binary mask, $\phi_{\tau}^{\text{lc}}$, is then used to modulate the backbone parameters, yielding the final policy $\pi_{\tau}$. Using the STE method, the agent updates the $\beta$ parameters and $\phi_{\tau}$ via backpropagation, keeping all the $\phi_{k}$ masks fixed.
Reward-Guided Initialization.
The trainable $\beta$ parameters are ideal to enable gradient descent to discover which policy is most beneficial on the current task. However, given the available information, one question is: What is a good initial weighting for the set of masks that include the local mask and those received from other agents? MOSAIC proposes a reward-guided weighting mechanism to set the initial values. While such values can be later tuned by gradient descent, an ablation study suggested that performing a good initial guess leads to a significant advantage. The intuition is that an agent with low performance may favor external knowledge, while an agent with higher returns could rely more on their own policies. The initial set of $\beta$ is given by
$\beta_{\tau}=0.5+0.5r,\quad\beta_{k}=\frac{0.5(1-r)}{|P_{c+1}|}\quad,$ (9)
where $r\in[0,1]$ is the agent’s normalized average return over the most recent iteration and $|P_{c+1}|$ is the number of peer masks received during a communication event. This scheme encourages a smooth shift from exploration to self-reliance as learning progresses.
The process just described occurs multiple times, and specifically, each time new masks are received. Therefore, before a new set of peer masks $P_{c+1}$ is integrated, the agent performs a consolidation operation that combines its current task mask $\phi_{\tau}$ with the set of previous peer masks $P_{c}$. This is done using a weighted linear combination similar to the forward pass. The resulting consolidated mask updates the existing $\phi_{\tau}$,
$\phi_{\tau}\leftarrow\beta_{\tau}\phi_{\tau}+\sum_{k=1}^{|P_{c}|}\beta_{k}\phi_{k}\quad.$ (10)
The consolidation reduces multiple currently held peer masks to one without affecting the policy, effectively maintaining scalability with respect to memory requirements.
2.5 Summary
The complete MOSAIC algorithm executes the processes described above in a loop. Figure 1 illustrates the algorithmic pipeline with each of the components. MOSAIC computes embeddings from batches of state-action-reward (SAR) data collected under the current policy, and these embeddings evolve over time as the policy improves. Knowledge reuse is governed by a linear combination of task-specific masks selected using similarity (computed from embeddings) and performance heuristics, with beta parameters re-initialized at each communication event and then fine-tuned to reflect downstream utility to the current task. While not part of this study, Appendix B provides a formalization of how masks can be linearly combined and how beta parameters can be initialized to extend MOSAIC to a fully-fledged lifelong reinforcement learning setting Ben-Iwhiwhu et al. (2023); Xie and Finn (2022).
Algorithm 1 provides an outline of the overall process. Pseudo-code for how the forward pass is computed is provided in Appendix E Algorithm 2. MOSAIC makes use of parallelized processes to achieve asynchronicity between learning and sharing knowledge. These parallel processes are described in Appendix E Algorithms 3, 4, and 5, which handle the updating of the mask and linear combination parameters on every communication event and manage the communication client/server, respectively. Queries and masks are shared around the system via multi-processing queues $Q_{\text{emb}}$ and $Q_{\text{mask}}$. Appendix E Algorithm 6 shows the computation of the Wasserstein task embeddings. The Python code of the complete algorithm to reproduce the results is provided at https://github.com/DMIU-ShELL/MOSAIC.
Algorithm 1 Agent Training Loop in MOSAIC
1:Assigned task $\tau\in\mathcal{T}$, shared backbone $\Phi$
2:Communication event interval $\Delta_{\text{comm}}$, embedding interval $\Delta_{\text{embed}}$
3:Replay buffer $\mathcal{D}_{\tau}$, rollout length $z$, total steps $\Omega$, cosine similarity threshold $\theta=0.5$, $N\leftarrow 128$
4:Initialize mask $\phi_{\tau}$, embedding $v_{\tau}\leftarrow\mathbf{0}$, return estimate $r\leftarrow 0$, step counter $t\leftarrow 0$
5:Initialize peer mask set $P_{0}=\emptyset$, and linear combination parameters ${\beta}\in\mathbb{R}^{1+|P_{0}|}$
6:Launch parallel threads:
7:    ReplacePeerMasks($Q_{\text{mask}}$) $\triangleright$ Receives and applies mask updates Alg. 3
8:    CommunicationModule($Q_{\text{emb}},Q_{\text{mask}}$) $\triangleright$ Handles queries and peer responses Alg. 4
9:while $t<\Omega$ do
10:     Roll out policy $\pi_{\tau}=\pi_{\Phi\odot g(\phi_{\tau}^{\text{lc}})}$ for $z$ $\triangleright$ Forward pass Alg. 2
11:     Store transitions in $\mathcal{D}_{\tau}$; update iteration return estimate $r$
12:     Update $\phi_{\tau}$ using policy gradient on loss $\mathcal{L}(\pi_{\tau})$triangleright Via Eq. 8
13:     $t\leftarrow t+z$
14:     if $t\bmod\Delta_{\text{embed}}=0$ and $|\mathcal{D}_{\tau}|\geq N$ then
15:         Sample $N$ samples from $\mathcal{D}_{\tau}$ and compute task embedding $v_{\tau}\leftarrow\psi(\mu_{\tau})$$\triangleright$ Alg. 6
16:     end if
17:     if $t\bmod\Delta_{\text{comm}}=0$ then
18:         Send TEQ for $(v_{\tau},r)$ to CommunicationModule via $Q_{\text{emb}}$ and acquire set of peer masks $P_{c+1}=\{\phi_{1}...\phi_{K}\}$ in CommunicationModule
19:         Update model with $P_{c+1}$ for forward pass via ReplacePeerMasks through $Q_{\text{mask}}$
20:     end if
21:     if $t=\Omega$ then
22:         Enter idle mode to keep CommunicationModule active
23:     end if
24:end while
3 Experiments
MOSAIC is evaluated on three sparse-reward reinforcement learning benchmarks: MiniHack MultiRoom (Samvelyan et al., 2021), MiniGrid (Chevalier-Boisvert et al., 2018), and an image sequence learning task from CT-graph (Soltoggio et al., 2019). Each benchmark presents a curriculum of tasks with underlying structural commonalities, but also significant differences that result in some tasks being unrelated to others. The objective is to assess whether MOSAIC can autonomously identify and exploit these relationships to outperform agents trained in isolation or agents that blindly exchange all knowledge. Performance graphs show averaged results and 95% confidence interval for various aggregations of experiments with a minimum of 5 seeds (Colas et al., 2018). The optimization of Eq. 1 employed PPO (Schulman et al., 2017) with a fully-connected network for the image sequence learning and MiniGrid tasks, and a convolutional network for MiniHack. The model architecture and hyperparameters (per module) are reported in the Appendix E (Figure 21 and Tables 4, 5, and 6).
3.1 Image Sequence Learning
This benchmark is an RL problem based on the Configurable Tree Graph (CT-Graph benchmark Soltoggio et al. (2019)), where a set of procedurally generated navigation tasks implement a variable length RL problem, and where each node in a tree-graph is an RL state encoded as image. We use this benchmark to define a curriculum of 28 tasks with navigation paths of increasing depths of the graph (2 to 8) from four independent sets of images. This setting creates four sets of unrelated tasks with seven related tasks in each set. The seven tasks in a set have an increasing depth of the navigation tree, generating a range of RL tasks from easy (depth 2) to very hard (depth 8), effectively creating a hierarchy of task’s complexity or difficulty. The sparse reward and exponential branching of the tree-graph structure make this benchmark a challenging RL setting, with reward probabilities as low as $\approx 7.74\times 10^{-9}$ for the most difficult tasks with depth 8 (Soltoggio et al., 2019).
Results in Figure 2(Top) show that MOSAIC achieves significantly faster learning and broader task coverage, outperforming MOSAIC-NoComm by 2.7$\times$, reaching a maximum total return of 26.0 across all 28 tasks. MOSAIC reaches 50% performance (14) in 37 iterations (18,944 steps). MOSAIC-NoComm plateaus at a maximum of 9.6. On average, MOSAIC-NoComm fails on 18 tasks, whereas MOSAIC fails on only 2.
3.2 MiniHack MultiRoom
MiniHack defines a grid-based navigation challenge with sparse rewards and pixel-based observations. Agents must navigate rooms and obstacles to reach a goal location. MOSAIC is evaluated on 14 navigation tasks organized into two clusters of increasing difficulty: one with room sizes of $4\times 4$, and another with $6\times 6$. Each level in a cluster adds an additional room, connected by a closed (but unlocked) door. The agent receives a reward of +1 for completing the task and a penalty of -0.01 for colliding with walls. The configuration of each task is randomly changed every episode, ensuring that agents must generalize across varying observations, goal locations, and door locations rather than memorize solutions.
In this experiment, agents must reuse sub-behaviors such as navigating to doors and exploring unseen regions to maximize performance across all tasks. Results in Figure 2(Bottom) show that MOSAIC agents collectively reach 0 return by iteration 132 (270,336 steps), compared to iteration 176 (360,448 steps) for MOSAIC-NoComm, which is a 25% reduction in the number of samples to reach zero reward. Over the full training run, MOSAIC achieves a maximum total return of 9.04 across the 14 tasks. MOSAIC-NoComm peaks at only 3.96.
Figure 2: Results comparing communicating MOSAIC agents versus isolated agents on the same set of tasks. (Top) Image sequence learning benchmark: average performance for 140 runs (28 tasks and five seeds/task) with 95% confidence intervals. (Bottom) MiniHack Multiroom benchmark: average performance for 70 runs (14 tasks and 5 seeds/task) with 95% confidence intervals. Sharing agents learn significantly better policies across more tasks in both benchmarks, with relative gains of 170.8% and 128.2% compared to the isolated learning baseline in the image sequence and MiniHack benchmarks, respectively.
3.3 MiniGrid Crossing
The MiniGrid benchmark presents a sparse-reward grid-world navigation problem with symbolic observations. MOSAIC is evaluated on 14 tasks that span seven SimpleCrossing variants and their seven LavaCrossing counterparts. Agents must navigate around obstacles or hazardous terrain and reach a goal location. Variations in these tasks arise from differences in the layout, placement of objects, and structure of the room. Agents must learn and reuse generalizable skills across different variations without compromising their policies.
This curriculum evaluates MOSAIC’s selection and composition mechanisms in a non-trivial task curriculum. In this setting, identifying reusable policies may be non-trivial, and tasks may have subtle interferences that may not be captured by the cosine similarity method. The evaluation focuses on whether knowledge acquired in SimpleCrossing tasks effectively transfers to LavaCrossing variants, as well as to other SimpleCrossing tasks. Figure 3 reports the summated return curves for each method. Table 1 summarizes both the final and intermediate performance metrics.
MOSAIC is compared against representative baselines. Multi-Task PPO (MTPPO) shares a backbone across all tasks without modularity. Multi-DQN (MDQN) (D’Eramo et al., 2020) attaches task-specific Q-heads to a shared encoder, enabling early selective feature reuse. Mixture of Orthogonal Experts (MOORE) (Hendawy et al., 2024) enforces expert diversity via Gram-Schmidt projection. PCGrad+MOE (Yu et al., 2020) mitigates gradient conflict using a fixed expert pool. These baselines span different approaches to parameter sharing, task specialization, and modularity. Results on the MiniGrid benchmark are shown in Figure 3.
Figure 3: Comparison of MOSAIC to baseline approaches on the 14 task MiniGrid curricula made up of SimpleCrossing and LavaCrossing task variations: average performance for 70 runs (14 tasks and five seeds/task) with 95% confidence intervals. In this experiment the reward function is given by 10.9(stepsmax $1-0.9(\frac{\text{steps}}{\text{max steps}})$ resulting in a total maximum of $\approx$12.
Table 1: Performance metrics on the MiniGrid Crossing benchmark. The table reports the final average return and the number of iterations to reach 25%, 50% and 75% of the theoretical maximum. Dashes indicate that the thresholds were not reached during the training.
MOSAIC MTPPO MDQN PCGrad+MoE MOORE MOSAIC-NC
Final perf 11.67 4.64 0.31 6.66 8.83 10.78
25% perf 55 48 - 24 13 46
50% perf 90 - - 54 24 87
75% perf 119 - - - - 291
MOSAIC outperforms all baselines in final return, achieving a maximum total return of 11.67. The performance gap ranges from 1.08× over MOSAIC-NoComm to 37.65× over MDQN. MOORE demonstrates stronger early-stage performance, reaching 25% and 50% of the total return in 13 (26,624 steps) and 24 (49,152 steps) iterations, respectively, approximately 76.4% and 73.3% faster than MOSAIC. PCGrad+MoE also displays faster improvement in earlier iterations, reaching 25% in 24 iterations (49,152 steps) and 50% in 54 iterations (110,592 steps). In MTRL systems, parameter sharing enables early gradient sharing from all tasks into a shared backbone network, which quickly improves shared features even from weak or random policies. In contrast, MOSAIC relies on modular mask composition that is driven by task embeddings. In earlier stages, each task-specific mask and embeddings reflect a random policy. Without initial reward signals, the linear combination of such masks is ineffective or even harmful. This results in a bottleneck until iteration 100, which the baselines avoid by implicitly averaging across tasks. Both MOORE and PCGrad+MoE fail to reach 75% performance, whereas MOSAIC continues improving, ultimately surpassing all baselines in both stability and final total return. MTPPO performs poorly, with a peak total return of 4.64, equating to 30% of the maximum theoretical performance. MDQN fails to learn any tasks in this particular benchmark, peaking at a total return of 0.31.
3.4 How Similarity Helps Targeted Policy Transfers
To provide a better insight on the utility of similarity in knowledge reuse, Figure 4 visualizes two key metrics: the average cosine similarity between task embeddings, and the average convergence values of the $\beta$ parameters, denoted $\hat{\beta}_{j}^{i}$, which are computed as average of the $\beta$ used in the consolidation operations (Eq. 10). Each $\hat{\beta}_{j}^{i}$ can be interpreted as a learned relationship from task $j$ to task $i$, estimating how useful a policy trained in task $j$ is for learning task $i$.
Figure 4: Visualization of pairwise cosine similarity and $\hat{\beta}$ statistics in the image sequence learning benchmark. (A) The cosine similarity matrix and (B) the $\hat{\beta}$ values are shown to visualize what tasks are similar and which policies are used for which tasks. (C) The cosine similarity values are clustered using WPGMA, and the results are illustrated in a dendrogram. (D) The cosine similarity matrix is replotted with tasks arranged according to the clustering. The four clusters of tasks emerge clearly. (E) The clustered $\hat{\beta}$ values show that similar tasks also have the highest indicator of policy reuse among each other. Annotated heatmaps of the individual clusters can be found in Figure 15 in the Appendix.
Figure 4(A) shows cosine similarities for tasks in a random order, which is a default situation for an arbitrary set of unknown tasks. These values are clustered using WPGMA in Figure 4(C), producing a dendrogram that now clearly highlights four main groups of similar tasks as well as possible hierarchical relationships. Reordering the matrix accordingly (Figure 4(D)) makes these clusters more visually salient. Figures 4(B) and (E) show the corresponding $\hat{\beta}_{j}^{i}$ matrices, before and after clustering. Notably, the clustered $\hat{\beta}$ have a similar pattern to the similarity values, indicating that similar tasks are also those that most heavily reuse each other’s policies during learning.
The tasks in the Image Sequence Learning benchmark were purposely devised to be organized into four distinct image sets, each containing seven tasks that vary only in trajectory depth, from shallow to deep. Within each group, cosine similarity scores tend to be higher between tasks that are close in depth and lower between those that are further apart. This pattern suggests that MOSAIC embeddings capture both the visual similarity induced by the image distribution and the structural similarity defined by specific optimal policies. While the similarity matrix is symmetrical by definition, the $\hat{\beta}$ values exhibit a more selective asymmetric reuse pattern that reflects the source-to-target nature of policy reuse. In addition, the initial weightings of $\beta$ are tuned to maximize performance. The analysis reveals that similarity and knowledge reuse are highly correlated.
Finally, we investigate how the reuse is affected by the progression of learning across multiple tasks and multiple agents. Figure 5 details the components of the performance graph (averaged in Figure 2(A)) by grouping tasks by the length (depth of the CT-graph) in the image sequence learning benchmark. The learning dynamics reveal an interesting emergent property of the system. The knowledge reuse is affected by the fact that easier tasks are learned early, and their policies enable agents solving progressively more complex tasks. This analysis also explains the 170.8% higher performance of communicating agents versus isolated agents in Figure 2(A), which derives from the crucial sharing processes that enable agents on difficult tasks to retreive, combine and refine policies from simpler but related tasks, and find solutions to the most complex problems.
Figure 5: Performance grouped by task complexity, i.e., seven different levels, corresponding to different graph’s depths, in the image sequence learning problem. Average performance for 20 runs (four tasks and five seeds/task) with 95% confidence intervals. (Top) Performance by difficulty level in MOSAIC. (Bottom) Performance by difficulty level in MOSAIC-NoComm, i.e., isolated agents. Tasks in MOSAIC are solved progressively, from the simplest to the hardest tasks. Isolated agents, on the other hand, only solve the two simplest tasks, and partially the third, but fail on the four most complex tasks.
3.5 Ablation Studies
Ablation studies were conducted on the image sequence learning problem (Section 3.1) to assess the effect on performance that emerges from the combination of the key algorithmic steps: the knowledge selection criteria and the weighted policy initialization. Figure 6 compares three ablated variants of MOSAIC: one without cosine similarity-based selection (criterion 1), one without performance-based selection (criterion 2), and one without reward-guided weight initialization. MOSAIC-NoComm is included as a reference baseline.
The ablation without cosine similarity-based selection (criterion 1) reaches a total return of 20.1. MOSAIC achieves 26.0, outperforming it by a factor of 1.29. We speculate that this gap could widen further in settings in which the number of unrelated tasks grows, i.e., where selecting the most relevant knowledge becomes critical.
The second ablation study without performance-based selection (criterion 2) performs comparably with MOSAIC at the end of the training (25.6), but it shows significantly slower learning. These dynamics make sense because, as all agents improve their performance on their respective tasks, selecting for high reward becomes less critical. These two ablation studies (removal of criterion 1 or 2) reach 50% of the theoretical maximum (14.0) at iterations 76 (38,912 steps) and 71 (36,352 steps), respectively. MOSAIC reaches this threshold at iteration 37 (18,944 steps), reaching 50% performance 1.9$\times$~2.0$\times$ faster than baselines.
The third ablation study removes the reward-guided initialization and instead uses a fixed weighting of 0.5 for the local policy and 0.5 for external policies. This experiment performs comparably with MOSAIC, however, the average return shows periodic instability, characterized by sharp performance drops coinciding with communication events, resembling a sawtooth pattern. This behavior suggests that the reward-guided initialization is a useful approach to maintain stability in performance when integrating external knowledge.
Further ablation studies are shown in Section C to test the impact of querying frequency on the performance of MOSAIC, and the impacts of the number of samples and size of the reference distribution on the Wasserstein task embeddings.
Figure 6: Ablation studies with average performance for 140 runs (28 tasks and five seeds/task) with 95% confidence intervals. The comparison shows the full MOSAIC algorithm and ablated variations: with no cosine similarity-based selection (criterion 1), no performance-based selection (criterion 2), and no reward-guided weight initialization. MOSAIC-NoComm is included for reference. The effectiveness of MOSAIC arises from the combination of all its components, the removal of which leads to a significant degradation in performance or learning behavior.
4 Discussion
The results, while showing different performance advantages across different tasks and curricula, indicate that the principles introduced in MOSAIC are effective in diverse settings. MOSAIC showed significantly faster and more successful learning dynamics with respect to isolated agents working on the same sets of tasks. These observations strongly suggest that knowledge composition enables effective reuse of previously learned policies across tasks and agents, supporting similar observations in a centralized single-agent lifelong learning setting (Ben-Iwhiwhu et al., 2023). In particular, masks from some tasks were observed to accelerate learning when selectively transferred to agents dealing with other tasks. We conclude that the selection and compositional mechanisms lead to the discovery of successful policies with a significantly better sample efficiency. We observed the advantages to be particularly significant in sparse-reward and hierarchical reinforcement learning environments (image sequence learning with varying depth), where agents learning difficult tasks benefit from acquiring policies from agents that are learning simpler tasks, supporting findings in Florensa et al. (2018); Nachum et al. (2018); Uchendu et al. (2023).
The ablation studies highlight the importance of the selection criteria–cosine similarity for relevance, reward-based selection for utility–as well as the reward-based weighting to optimize the initial policies. The results of the ablation studies are particularly significant because they suggest that knowledge sharing and integration may not be effective in a vanilla setup where all knowledge is shared (Isele and Cosgun, 2018). It is conceivable that such criteria are crucial for scalability, for example, when implementing a system of a large number of interconnected agentic AI working on a large set of diverse tasks in a fully asynchronous setting (Foerster et al., 2016).
Interestingly, removing the reward-guided initialization of policies introduces periodic instability. Despite achieving a slightly higher final return (26.4 vs. 26.0), this variant suffers sharp drops after communication events, suggesting that newly acquired policies can interfere with learning when not weighted adaptively. It is also conceivable that this setting could become critical for the safe deployment of MOSAIC: new externally acquired policies, even if vetted by criteria 1 and 2 that should ensure relevance and performance, may nevertheless require “careful” introduction in the agent’s own policy (Amodei et al., 2016).
MOSAIC currently relies on raw episodic reward as a selection and weighting signal, which may limit generalization across environments with unstandardized or sparse reward functions. Future variants could use reward-normalized scores or task-progress metrics to support more environment-agnostic policy reuse (Jaques et al., 2019). Additionally, MOSAIC assumes a fully connected communication topology, which may not scale well in systems with many agents or bandwidth constraints; exploring adaptive or sparse topologies could improve scalability to real-world applications (Kaimakamidis et al., 2024; Marino et al., 2025; Tang et al., 2024). The current design also restricts mask composition to positive, additive combinations, limiting the ability to suppress interference from poorly aligned peer policies. Introducing signed or non-linear composition strategies could address this issue (He et al., 2022). Finally, the use of a shared, frozen backbone across agents may constrain representational diversity and reduce adaptability in highly heterogeneous task settings. Allowing for backbone specialization or dynamic composition techniques may improve transfer and robustness in more diverse environments (Houlsby et al., 2019).
While this work uses masking as the modular knowledge representation, similar mechanisms based on parameter isolation could be tested. For example, LoRA-like modules or adapters have shown significant promise in transformer architectures and effectively act as pre-trained modules that have been shown to work also with various fusion and combination approaches (He et al., 2022; Wu et al., 2024; Han et al., 2021; Qiu et al., 2020). This consideration opens the possibility for compositional reuse in transformer-based models and large-scale language agents.
Communication and selection in MOSAIC were introduced as a proof-of-concept to assess the validity of the hypothesis that knowledge selection (what and when to transfer (DARPA/MTO, 2021; Soltoggio et al., 2024)) is a key element in agentic, decentralized, and fully asynchronous systems. To achieve an even higher level of agency in MOSAIC agents, more aspects related to communication could be learned by a policy (Foerster et al., 2016). Agents learning and optimizing their communication could lead to more dynamic behaviors that are better suited to each agent’s task(s) and resources. However, learnable and dynamics communication policies could lead to potential safety concerns, as agents that learn to collaborate could learn greedy, or uncooperative policies that may benefit the agent but not the collective as a whole. In the simplest instance, an agent that does not share knowledge exploits the collective knowledge without contributing to it. In a worst-case scenario, malicious agents could share harmful policies. Such an extension of MOSAIC could lead to complex learning and collaboration dynamics described in Game Theory (Myerson, 2013). In summary, we are aware that the MOSAIC algorithm raises safety considerations (Hammond et al., 2025; Bengio et al., 2025): without appropriate filters or constraints, agents may transfer misaligned or flawed policies, leading to instability or harmful behavior in sensitive environments. MOSAIC-like systems are also vulnerable to attacks by malicious actors who exploit the peer-to-peer nature, poisoning models, or extracting policies or training data (Vassilev et al., 2024).
The knowledge composition mechanism in MOSAIC derives from approaches in lifelong machine learning, where agents encounter a stream of continuously changing tasks. Therefore, the algorithm is particularly suitable to be extended to a lifelong learning setting. In such an extension, each policy is a composition of three sources: a task-specific mask for the current task, prior masks from previously learned tasks, and peer-acquired masks from other agents. Effectively, MOSAIC is ideally designed to reuse prior knowledge from previous tasks when learning a continuous stream of tasks (Serra et al., 2018). For these considerations, we hypothesize that the reuse and integration of policies acquired from other agents within a set of own-policies could enable scalable and open-ended lifelong learning (Leibo et al., 2019; Stanley, 2019) across diverse task streams.
5 Conclusion
This paper introduced MOSAIC, a collaborative agentic algorithm for policy sharing and composition in reinforcement learning. By exploiting network masks with reward- and cosine similarity-driven selection, MOSAIC enables effective and efficient knowledge sharing, producing a collaborative collective system whose performance is significantly superior than that of isolated agents. The findings presented in this study support the broader hypothesis that effective knowledge transfer and reuse is possible in agentic systems supported by appropriate knowledge selection and integration principles. Interestingly, the proposed algorithm improves learning efficiency and effectiveness by exploiting parallel agentic learning that is shown to result in naturally emergent curricula, where agents learning easy tasks achieve their goals sooner, and subsequently support agents learning harder tasks. The capabilities shown by MOSAIC provide a proof-of-concept for a collaborative and fully autonomous AI collective composed of independent and cooperating agents.
Broader Impacts
MOSAIC introduces a collaborative and autonomous knowledge-sharing algorithm to enable agentic systems to seek and share machine-learned knowledge. Potential positive impacts include advancing autonomous learning in real-world multi-agent settings, such as agentic language models, robotics, disaster response, multi-agent search & rescue robotic teams, satellite swarms, or systems where privacy, data ownership, legislation, or bandwidth constraints preclude centralized coordination. The policy composition approach in MOSAIC promotes interpretability, as task compositions can be inspected and traced back to specific prior knowledge sources. Policy reuse may extend across domains of language, perception, and control, enabling more transparent and collaborative AI systems in human-facing applications.
The autonomous knowledge-sharing capabilities of MOSAIC also pose new risks. Agents could transfer flawed or misaligned policies at a fast rate, effectively spreading potential malicious information across a large network, a process that could be difficult to reverse. Future work should focus on developing mechanisms to ensure policy transfers are safe and aligned with each agent’s objectives.
Acknowledgments
This material is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under contract No. HR001121901 (Shared Experience Lifelong Learning) and the Industrial Robots-as-a-Service (IRaaS) project funded by the EPSRC (EP/V050966/1).
Appendix A Related Work
The section provides additional information on the related work that contributed to MOSAIC.
Modular Representations and Masking.
Modular policies promote parameter reuse by isolating task-specific substructures. Binary masks define sparse sub-regions of shared networks, reducing interference and enabling structured composition [Mallya et al., 2018, Mallya and Lazebnik, 2018, Ramanujan et al., 2020, Serra et al., 2018, Wortsman et al., 2020, Koster et al., 2022]. Some approaches compose a mask to support task transfer [Ostapenko et al., 2021, Mendez et al., 2022, Gummadi et al., 2022, Ben-Iwhiwhu et al., 2023]. Mixture-of-Experts (MoE) architectures route inputs to sparse expert modules via learned gates [Shazeer et al., 2017, Fedus et al., 2022], with variants adapted to multi-task and lifelong settings [Gururangan et al., 2020, Mendez and Eaton, 2021]. MOORE imposes orthogonality constraints across experts to promote representational diversity [Hendawy et al., 2024]. Related work in modular meta-learning emphasizes structural reuse through task-level module composition [Alet et al., 2018]. LoRA adds task-specific low-rank adapters to fixed models for efficient specialization [Hu et al., 2022]. Neural Module Networks support compositional inference via dynamically assembled submodules [Andreas et al., 2016]. MultiCriticAL isolates value functions per task to mitigate interference in actor-critic training [Mysore et al., 2022]. Recent robotic lifelong learning systems also adopt modular structures with Bayesian non-parametric inference to preserve and combine task-specific knowledge [Meng et al., 2025].
Distributed Knowledge Reuse.
Modular reuse has been extended to distributed and federated RL. FedWeIT [Yoon et al., 2021] and CONFEDMADE [Paul et al., 2025] use attention-weighted masks or modular encoders to isolate transferable components. Other methods support cross-agent sharing of task-specific masks or subnetworks [Gao et al., 2021, Ge et al., 2023, Nath et al., 2023, Rostami et al., 2017, Gerstgrasser et al., 2023]. Recently, PRISM [Seo et al., 2025] introduced a federated generative framework that communicates stochastic binary masks instead of full parameters. FedProto [Tan et al., 2022] facilitates communication-efficient reuse via prototype aggregation. DiPaCo [Douillard et al., 2024] and OpenDILoCo [Jaghouar et al., 2024] offer decentralized, low-communication frameworks for collaborative model training. Recent perspectives such as Soltoggio et al. [2024] articulate a broader vision of decentralized collective AI, emphasizing modular architectures that enable asynchronous, interpretable knowledge exchange at scale. Diskin et al. [2021], McMahan et al. [2017], and Meng et al. [2025] also explore communication-efficient knowledge preservation and reuse in lifelong learning agents.
Peer-to-peer frameworks exchange representations via role-switching, consensus updates, or parameter interpolation [Sengupta et al., 2024, Alvar and Bajić, 2019, Kaimakamidis et al., 2024, Yu et al., 2022, Nadiger et al., 2019, Lozano-Cuadra et al., 2025]. PaCo and similar methods interpolate within shared subspaces to define composable policies [Sun et al., 2022]. Model-based frameworks support decentralized planning with shared structured representations [Jiang et al., 2021, Du et al., 2025, Shi et al., 2022]. PEEPLL introduces confidence-based peer communication for decentralized lifelong learning, reducing reliance on environmental supervision by leveraging self-evaluated knowledge sharing between agents [Tarale et al., 2025]. CenRA [Ma et al., 2025] centralizes reward broadcasting, while jump-starting methods guide early-stage learning via reuse [Uchendu et al., 2023].
Task Similarity and Inference.
Task inference enables modular adaptation without task labels. Latent and Bayesian models estimate embeddings from trajectories [Zintgraf et al., 2021, Achille et al., 2019], while non-parametric and prototype-based methods define task similarity via behavioral statistics [Yarats et al., 2021, Chu et al., 2024, Liu et al., 2022, Kolouri et al., 2021, Dick et al., 2024]. Wasserstein barycenter methods enable structure-aware ensembling by incorporating semantic side information through optimal transport distances [Dognin et al., 2019], and learning embeddings into entropic Wasserstein spaces offers a flexible alternative to Euclidean embeddings for capturing rich relational structures [Frogner et al., 2019]. Contrastive objectives such as CURL [Laskin et al., 2020] and ATC [Stooke et al., 2021] improve representation learning, though they typically require encoder training and large batch sizes. Structure-aware distillation has also been explored in incremental graph-based settings [Wang et al., 2021], offering another avenue for modular task inference.
Gradient-based meta-RL methods such as MAML [Finn et al., 2017] optimize for rapid task adaptation through fine-tuning, while variational approaches like VariBAD [Zintgraf et al., 2021] condition policy learning on inferred latent variables. In contrast, non-parametric methods that reuse task-conditional modules based on similarity measures offer an alternative path to fast generalization without the need for gradient-based adaptation.
Model Merging and Upcycling.
Emerging techniques enable the combination and repurposing of existing models through weight-space merging or routing. Wortsman et al. [2022] and Ramé et al. [2023] propose interpolating parameters from multiple fine-tuned models to improve generalization. TIES-Merging [Yadav et al., 2023] resolves interference between merged models. Komatsuzaki et al. [2023] and Sukhbaatar et al. [2024] demonstrate sparse upcycling: extracting sparse modular experts from dense checkpoints. These strategies provide alternatives to training from scratch and are critical for decentralized reuse.
Adaptive and Elastic Architectures.
Recent work explores modular and resource-aware architectures. MatFormer [Devvrit et al., 2024] and Mixture-of-Depths [Raposo et al., 2024] dynamically allocate compute and adjust routing depths in transformer models. These approaches support efficient inference and specialization while preserving modular structure, principles aligned with our use of task-conditional module composition.
Hierarchical and Curriculum Learning.
Hierarchical RL tackles long-horizon credit assignment and sparse reward signals. HAC [Levy et al., 2019] introduces layered actor-critic policies operating at different temporal resolutions. CAMRL [Huang et al., 2023] proposes curriculum-driven asymmetric MTRL between teacher and student agents. Our method complements these paradigms by achieving curriculum-like generalization via unsupervised trajectory similarity and decentralized reuse.
Appendix B Extending MOSAIC to Lifelong Reinforcement Learning
MOSAIC can be naturally extended to a lifelong learning (LL) setting [De Lange et al., 2022, Kudithipudi et al., 2022], in which each agent learns a sequence of tasks while minimizing catastrophic forgetting and promoting forward transfer. LL MOSAIC agents can exploit their knowledge integration capabilities to accumulate knowledge that may derive from other agents, or from their own learning on previously encountered tasks. MOSAIC employs knowledge composition approaches that were developed in the field of lifelong reinforcement learning [Ben-Iwhiwhu et al., 2023]. In the experimental settings for MOSAIC, however, we employed a one task per agent approach to maintain the focus on knowledge sharing and composition. This setting can be extended to a more general case in which one agent is not limited to learning one task. Therefore, MOSAIC can be easily adapted to a fully-fledged lifelong learning scenario by modifying the linear combination and $\beta$ parameter adjustment.
Let agent $i$ currently be learning task $\tau_{k}+1$. Its mask is composed from three sources: the current task mask $\phi_{\tau_{k}+1}$, prior task masks $\{\phi_{\tau_{k}}\}_{k=1}^{K}$, and peer masks $\{\phi_{j}\}_{j=1}^{n}$, acquired from other agents. The composed mask for task $\tau_{k+1}$ for the forward pass is thus given by,
$\phi_{\tau_{k+1}}^{\text{lc}}=\beta_{\tau_{k+1}}\phi_{\tau_{k+1}}+\sum_{k=1}^{k}\beta_{\tau_{k}}\phi_{\tau_{k}}+\sum_{j=1}^{n}\beta_{j}\phi_{j},$ (11)
where each $\beta$ is a scalar weight normalized via softmax. In this case, the initialization problem includes setting the weighting for previous task masks, the current task mask, and the newly acquired masks from other agents. One simple approach could simply be to extend Eq. 9 to
$\beta_{\tau_{k+1}}=\frac{1}{3}+\frac{2}{3}r,\quad\beta_{\tau_{k}}=\frac{(1-r)}{3k},\quad\beta_{j}=\frac{(1-r)}{3n}.$ (12)
The addition of lifelong learning presents an additional challenge. Boundaries in the sequence of tasks learned by an agent need to be identified, so that a new mask can be learned for each new task. One solution to this could be to use the cosine similarity computed from the Wasserstein embeddings to monitor the incoming data stream and decide when a new task starts using statistical methods [Dick et al., 2020, 2024].
Appendix C Additional Ablation Studies
Additional ablation studies An additional ablation study was conducted to assess the effect of communication frequency on MOSAIC’s performance in the image sequence learning benchmark. Specifically comparing frequencies of 1, 5, 10 (default), 25, and 40 iterations, where the frequency indicates how often agents query and integrate peer masks. As shown in Figure 7, overly frequent communication (i.e., frequency of 1) results in performance degradation. In this setting, newly acquired peer masks continually override the agent’s own beta parameter adjustments, preventing meaningful adaptation and fine-tuning outside of the performance-based adjustment. Frequencies of 5 and 10 yield comparable early performance, with a frequency of 10 achieving higher final return. Conversely, infrequent communication (25 or 40 iterations) reduces the opportunity for useful knowledge transfer, leading to weaker overall performance. These results highlight the importance of balancing between querying for peer knowledge often enough for a benefit to occur, and allowing agents to then tailor peer masks towards their own tasks.
Figure 7: Comparison of MOSAIC with varying communication frequencies highlights the effect on performance in the image sequence learning problem. Striking the right balance in communication frequency is critical: infrequent communication can lead to missed performance gains, while communicating at every iteration may not allow sufficient time for effective local adaptation.
Figures 8 and 9 show results for a grid-search experiment on the number of samples $N$ used to compute an embedding (Equation 3) and the reference size $M$ (Equation 4). Heatmaps show the cosine similarities (Figure 8) and beta parameters (Figure 9) for each source-target task pair in the image sequence learning problem. In this experiment $N$=128 and $M$=50 denote the default configuration of MOSAIC in all experiments in this paper. The results indicate minimal differences as the value of $M$ increases from 50. There is an increase in the number of masks shared between tasks of different distributions as $M$ decreases.
Table 2 reports the proportion of sharing within vs. cross-task distribution, for average cosine similarity and average beta parameter values. Within-cluster cosine similarity remains high across all configurations (0.84–0.88), indicating robust internal coherence among task groups. The cosine similarity between task distributions is more variable (0.38–0.43), with higher values observed at lower $M$, suggesting decreased separability between groups when mask capacity is restricted. Beta parameters exhibit a strong asymmetry: cross-distribution reuse is negligible (~0.000–0.003), while reuse within each cluster remains stable (~0.036–0.041). These trends confirm that agents primarily leverage intra-cluster task knowledge and that both sampling and masking hyperparameters influence the clarity and utility of the learned compositional structure.
Table 2: The average cross-cluster and within-cluster cosine similarity and beta parameter values across various $N,M$ configurations. Average cosine similarity within clusters decreases as $N$ and $M$ increase. The average beta parameter values also increase as $N$ and $M$ increase, suggesting increased transfer between tasks driven by more accurate cosine similarity.
Metric N32 M50 N64 M50 N128 M10 N128 M20
Cross-Similarity 0.3887 0.3895 0.4288 0.4135
Within-Similarity 0.8779 0.8677 0.8425 0.8482
Cross-Beta 0.0021 0.0006 0.0029 0.0019
Within-Beta 0.0384 0.0368 0.0395 0.0377
Metric N128 M30 N128 M50 N128 M70 N128 M100
Cross-Similarity 0.3993 0.4047 0.4054 0.3956
Within-Similarity 0.8559 0.8529 0.8576 0.8571
Cross-Beta 0.0013 0.0006 0.0006 0.0000
Within-Beta 0.0374 0.0386 0.0397 0.0411
Figure 8: Cosine similarity heatmaps for configurations of $N$ and $M$. Cross-cluster cosine similarity increases with smaller values of $N$ and $M$, leading to instances of cross-cluster sharing. Higher values increase the accuracy of embedding cosine similarity.
Figure 9: Beta parameters heatmaps for configurations of $N$ and $M$. Poorer cosine similarity estimates result in the transfer of masks. This then leads to the reuse of these masks in the linear combination.
Appendix D Additional Experiment Analysis
Figure 10 illustrates an example of the sequence-learning problem based on [Soltoggio et al., 2019]. The agent must navigate a series of nodes labeled as H (home), W (wait state), D (decision state), and R (reward state). For the experiments in this paper, the wait state probability is set to zero, resulting in reduced complexity.
Figure 10: Illustration of the image sequence learning problem. Each graph is comprised of various states, represented by procedurally generated images. The agent must learn to navigate these image states and acquire a sparse reward. (Top) shows an example of a depth-1 graph. Bottom shows a depth 11 graph along the same trajectory.
Figure 11 shows examples of the MiniGrid curriculum used in our MiniGrid Crossing experiments. Agents are expected to learn useful behaviors/trajectories to navigate around obstacles in the form of walls or lava, and reach the goal state (green). This setup tests to see if agents can identify similarities in the structure of the tasks, irrespective of the changing observations (wall, lava), and transfer useful policies from SimpleCrossing to LavaCrossing variations, as well as on other SimpleCrossing/LavaCrossing variations.
Figure 11: Illustrations of the MiniGrid SimpleCrossing and LavaCrossing curriculum used in our experiments. Each SimpleCrossing variation is accompanied by a LavaCrossing counterpart. This approach tests whether MOSAIC agents can transfer knowledge across different observation spaces while maintaining agent specialization.
Figure 12 shows an example of a MiniHack MultiRoom task. The start and goal states are represented by squares with up and down arrows, respectively. The agent is expected to navigate through each room, learning to navigate towards and open doors along the way until it reaches the goal state in the final room.
Figure 12: Example illustration of the MiniHack MultiRoom benchmark. Each environment consists of multiple rooms of varying sizes, connected by closed (but unlocked) doors. Tiles with upwards arrows indicate starting locations, and tiles with downwards arrows indicate goal locations.
Figures 5, 13, and 14 show the individual task performance curves for the image sequence learning setting, MiniGrid Crossing, and MiniHack MultiRoom, respectively.
Figure 13: Individual task performance on the MiniGrid benchmark. (Top) MOSAIC: average performance on five seeds and 95% confidence intervals. (Bottom) MOSAIC-NoComm.
Figure 14: Individual task performance on the MiniHack benchmark. (Top) MOSAIC: average performance on five seeds and 95% confidence intervals. (Bottom) MOSAIC-NoComm.
D.1 Further Analysis of Similarity and Converged Beta Parameters
The image sequence learning benchmark consists of four distinct image datasets, resulting in four independent state spaces. Four unique trajectories are chosen, one for each image dataset. Seven hierarchically related tasks are selected from each trajectory-dataset combination resulting in a total of 28 tasks with four clusters denoted by (0 to 6), (7 to 13), (14 to 20), (21 to 27). Figure 15 shows the annotated results of the heatmaps presented in Figure 4(D) and (E) for each of these clusters, identified by the WPGMA (Figure 4(C)). In Figure 16, each pair of tasks is categorized as "Within Distribution" (e.g., (0-1) and "Across Distributions" (e.g., 0-14).
In the MiniGrid benchmark, tasks are divided into two families: SimpleCrossing (SC; tasks 0–6) and LavaCrossing (LC; tasks 7–13). Each SC task has a corresponding LC task with a matched structural layout but different obstacles (walls versus lava), forming seven pairs of canonical tasks (e.g. 0-7, 1-8) that are structurally matched. Surprisingly, the pairwise cosine similarity heatmap reveals that these canonical pairs do not have the highest similarity (see also Figure 18). Instead, many within-family pairs show stronger similarity, particularly among SC tasks. Many non-structurally matched pairs also exhibit unexpectedly high similarity. These results suggest that the Wasserstein task embeddings are able to encode non-trivial task relationships which MOSAIC can then leverage for knowledge selection.
The corresponding $\hat{\beta}$ matrix indicates that knowledge reuse in MiniGrid is selective and is not determined solely by the canonical structure. Reuse tends to favor structurally simple or early-learned tasks rather than those with the highest embedding similarity. Figure 18 confirms this pattern: canonical pairs show lower average similarity than both within-group and non-structurally-matched pairs, suggesting that MOSAIC’s embedding space captures aspects of task dynamics that go beyond explicit structural symmetry. Reuse decisions are shaped not only by similarity but also by policy quality and learning progress. This pattern may also reflect the temporal bias in reuse. Tasks such as SC0, which reach high performance early in training, become preferred sources of knowledge. Structurally matched tasks that appear later in training are often underutilized, due to the performance-based selection mechanism (criterion 2) that filters knowledge transfer based on relative performance.
The MiniHack benchmark extends the MiniGrid MultiRoom curriculum. It consists of tasks ranging from two to eight rooms in size, with task IDs 0–6 corresponding to room size $4\times 4$ and tasks 7–13 to size $6\times 6$. Canonical pairs are defined by room count; for example, task 0 and task 7 are both 2-room versions of different sizes. Cosine similarity analysis reveals clear grouping by room size, with tasks within the same spatial resolution exhibiting stronger similarity than canonical pairs across sizes, indicating that visual and navigational complexity induced by room dimensions plays a more significant role in shaping the embedding space than shared structural depth alone.
The matrix $\hat{\beta}$ shows that knowledge reuse in MiniHack occurs predominantly within the same room size group. Smaller tasks, such as the 2-room and 3-room configurations, are reused more frequently and earlier in training, consistent with their faster convergence. As in MiniGrid, structural pairing alone does not guarantee high reuse. Figure 20 shows that canonical pairs (e.g., 0–7, 1–8) have only moderate similarity, while many within-group pairs yield stronger transfer potential. The dominance of reuse within room size groups may also be explained in part by the timing of performance gains. Tasks with smaller rooms converge earlier and thus become eligible for reuse earlier, which reinforces their influence even when structurally similar tasks emerge later in training.
Figure 15: Individual heatmaps of each task distribution cluster from the image sequence learning results shown in Figure 4. Cosine similarity is highest between each task and the next task. Beta parameters indicate that cosine similarity has an inversely proportional relationship with cosine similarity, with foundational prior tasks having more weight. E.g., task 0 has the most weight for reuse in task 5.
Figure 16: Violin plot showing the distribution of pairwise cosine similarities between task embeddings in the image sequence learning benchmark. Pairs are grouped as either Within Distribution, where both tasks share the same image dataset and trajectory (e.g., Dataset 1-Dataset 1), or Across Distributions, where tasks come from different datasets and trajectories (e.g., Dataset 1–Dataset 3). Tasks within the same distribution show significantly higher similarity, indicating that MOSAIC embeddings reflect shared perceptual features and curriculum alignment.
Figure 17: Pairwise (Left) cosine similarity and (Right) converged beta parameters results from the MiniGrid experiment. Results are clustered using the WPGMA method, revealing two distinct distributions. Task IDs 0-6 correspond to the SimpleCrossing. Tasks 7-13 correspond to the LavaCrossing. The results indicate that the cosine similarity of Wasserstein task embeddings is able to uncover meaningful relationships.
Figure 18: Average cosine similarity between task embeddings grouped by structural and dynamic task relationships. The plot categorizes task pairs into three groups: Structurally Matched Pairs (SC–LC), consisting of canonical SimpleCrossing–LavaCrossing pairs with identical map layouts; Same Task Type (SC–SC or LC–LC), where both tasks belong to the same dynamics family but differ in layout; and Non-Structurally Matched Pairs, which includes all remaining SC–LC task pairs not matched by structure. Structurally matched pairs do not show higher similarity than within-type pairs, suggesting that embedding similarity is more influenced by task dynamics than map layout.
Figure 19: Pairwise (Left) cosine similarity and (Right) converged beta parameters results from the MiniHack experiment. Results are clustered using the WPGMA method.
Figure 20: Average cosine similarity between task embeddings grouped by room structure and scale. Task pairs are grouped into three categories: Same Room Size, where both tasks have the same room size (either 4×4 or 6×6); Same Room Count (Different Size), which refers to canonical pairs having the same number of rooms but different room sizes; and Different Room Size and Count, which includes all remaining task pairs differing in both dimensions. The results show that spatial scale exerts a dominant influence on embedding similarity, while topological alignment across scales yields no increase in similarity.
Appendix E Implementation Details
Algorithm 2 describes how the forward pass is computed using a weighted linear combination of the current task mask, peer masks, and the backbone network. Algorithm 5 presents an alternative communication strategy in MOSAIC, where agents perform selection on the server side by returning only the most relevant masks from their knowledge base to querying agents. This alternative approach improves scalability and communication efficiency without changing the way MOSAIC works. This alternative approach is especially useful in the lifelong learning scenario where each agent encounters multiple tasks. By reducing the exchange of irrelevant information, this server-side selection mechanism enables MOSAIC to adapt effectively to lifelong learning, where agents would be expected to learn multiple task-masks. Rather than share all information of all the available knowledge, peer MOSAIC agents would perform selection locally for the querying agents and send the appropriate mask(s). The lifelong learning adaptation of MOSAIC is discussed in Section B.
Algorithm 2 Forward pass using masks for network layer $l$
1:Masks $\phi_{\tau}$, peer mask set $P=\{\phi_{k}\}_{k=1}^{|P|}$, parameters $\beta_{\tau}$, $\{\beta_{k}\}_{k=1}^{|P|}$, backbone $\Phi$
2:procedure ForwardPass(x)
3:     if $|P|=0$ then $\triangleright$ no peer masks available
4:         Set $\phi_{\tau}^{\text{lc}}\leftarrow\phi_{\tau}$
5:     else
6:         $\phi_{\tau}^{\text{lc}}\leftarrow g\left(\beta_{\tau}\phi_{\tau}+\sum_{k\in P}\beta_{k}\phi_{k}\right)$triangleright Eq. 8
7:     end if
8:     $\pi_{\tau}\leftarrow\pi_{\Phi\odot\phi_{\tau}^{\text{lc}}}$$\triangleright$ Modulate backbone with LC mask to get policy
9:     $y\leftarrow\pi_{\tau}\cdot\textbf{x}$$\triangleright$ Compute output from policy
10:     return $y$
11:end procedure
Algorithm 3 Updating set of peer masks and resetting linear combination parameters
1:latest iteration return $r$
2:procedure ReplacePeerMasks($Q_{\text{mask}}$)
3:     while running do
4:         Wait for set of new peer masks $P_{c+1}=\{\phi^{\text{new}}_{1},\dots,\phi^{\text{new}}_{K}\}$ from $Q_{\text{mask}}$
5:         Update $\phi_{\tau}\leftarrow\beta_{\tau}\phi_{\tau}+\sum_{k\in P_{c}}\beta_{k}\phi_{k}$$\triangleright$ Consolidation step
6:         Cache consolidated mask for forward pass
7:         Update $\beta_{\tau}\leftarrow 0.5+0.5r$, $\beta_{k}\leftarrow\frac{0.5(1-r)}{|P_{c+1}|}\quad\forall k\in P_{c+1}$
8:         $P_{c}\leftarrow P_{c+1}$
9:     end while
10:end procedure
Algorithm 4 MOSAIC Communication Module
1:Agent’s IP/Port $\texttt{addr}_{i}$, Communication graph $\mathcal{G}=\{\mathcal{I},\mathcal{E}\}$, where =$\mathcal{E}=\{(i,j)\mid i,j\in\mathcal{I},\ i\neq j\}$
2:procedure CommunicationModule($Q_{\text{emb}},Q_{\text{mask}}$)
3:     Initialize parallel process CommunicationServer($Q_{\text{mask}}$)
4:     while running do
5:         Wait for query $(v_{i},r_{i},\texttt{addr}_{i})$ from $Q_{\text{emb}}$
6:         Send query to known peers
7:         Wait for set of peer QRs $\{(v_{j},r_{j},\text{mask\_id}_{j},\texttt{addr}_{j})\}$ from CommunicationServer
8:         Select peers $P\leftarrow\{j\ |\ \mathbb{I}_{\text{align}}\text{ and }\mathbb{I}_{\text{perf}}\}$
9:         Send mask request (MR) to selected peers $P$
10:         Wait for peer masks from CommunicationServer and send to model via $Q_{\text{mask}}$
11:     end while
12:end procedure
Algorithm 5 Communication server
1:World size $W$, Agent’s IP/Port $\texttt{addr}_{i}$
2:procedure CommunicationServer($Q_{\text{mask}}$)
3:     Bind TCP socket to $(\texttt{addr}_{i})$
4:     Listen for connections with backlog size $W$
5:     while running do
[…]
7:         Receive message from connection
8:         if message is an TEQ then
9:              if Server-side selection then
10:                  Select mask(s) and respond with MTR to remote_addr
11:              else
12:                  Respond to remote_addr with embedding and performance (QR)
13:              end if
14:         else if message is an QR then
15:              Store QR $(v_{j},r_{j},\text{mask\_id},\texttt{addr}_{j})$ in set of query responses
16:         else if message is an MR then
17:              Fetch corresponding mask for mask id and respond to remote_addr with MTR
18:         else if message is an MTR then
19:              Add mask to set of acquired peer masks
20:         end if
21:     end while
22:end procedure
Algorithm 6 Wasserstein Embedding Computation
1:procedure ComputeEmbedding($\mathcal{D}$)
2:     Compute embedding $v_{\tau}^{\text{new}}\leftarrow\psi(\mathcal{D})$triangleright via Eq. 6
3:     $v_{\tau}\leftarrow\frac{1}{2}(v_{\tau}+v_{\tau}^{\text{new}})$triangleright Moving average update
4:     return $v_{\tau}$
5:end procedure
Tables 4, 5, and 6 report all hyperparameters used for the PPO model, communication module, and embedding module, respectively. Figure 21 reports details on the model architectures used in our experiments. All seeds used in our experiments can be found in the code implementation: https://github.com/DMIU-ShELL/MOSAIC. Table 3 lists the hardware devices used in the experiments.
Figure 21: Details on architectures used for our experiments. (Left) shows the fully linear model used in the image sequence learning and MiniGrid benchmarks. (Right) shows the convolutional model used for the MiniHack benchmark. All models use ReLU activation. Blue depicts pass-through layers (these are not masked), green depicts linear masking layers, and pink depicts convolutional masking layers.
Table 3: Compute resources used to run the experiments presented in this paper for MOSAIC and baselines. We made use of the MiG configurations that are supported by the NVIDIA A100 GPUs to run each MOSAIC agent with its own instance. This configuration limits the effects of one agent on another from a hardware perspective.
System 1
CPU AMD EPYC 9634 (168 logical cores)
GPU(s) 2× NVIDIA A100 80GB PCIe
GPU Driver 535.216.03
CUDA Version 12.2
RAM 512 GB
System 2
CPU AMD EPYC 7713P (128 logical cores)
GPU(s) 2× NVIDIA A100 40GB PCIe
GPU Driver 535.216.03
CUDA Version 12.2
RAM 256 GB
System 3
CPU AMD Ryzen 9 5950X (32 logical cores)
GPU(s) 1× NVIDIA GeForce RTX 4090
GPU Driver 535.183.01
CUDA Version 12.2
RAM 128 GB
Table 4: PPO and model hyperparameters used in our experiments across each benchmark.
Hyperparameter Image Sequence Learning MiniGrid MiniHack
Learning rate 0.00025 0.00025 0.00025
Backbone seed 9157 9157 9157
Num. workers 1 1 1
Discount factor 0.99 0.99 0.99
GAE Tau 0.99 0.99 0.99
Entropy 0.01 0.01 0.01
Rollout length 512 2048 2048
Opt. iterations 8 8 8
Num. mini batches 64 64 64
PPO ratio clip 0.1 0.1 0.1
Gradient clip 5 5 5
Total training steps 102400 1024000 2048000
Table 5: Hyperparameters specific to the communication module of the system. Querying frequency refers to the frequency of queries with respect to which iteration the agent is on. Query wait and Mask wait refer to the built-in wait delays in the peer-to-peer communication protocol. One of the primary goals of MOSAIC is asynchronicity, which means agents cannot afford to wait for extended periods for information to arrive. We found that 0.3 ms worked well in our experiments, though this may not work well when scaled to more agents.
Hyperparameter Image Sequence Learning MiniGrid MiniHack
Query freq. (iterations) 10 25 10
Query wait (ms) 0.3 0.3 0.3
Mask wait (ms) 0.3 0.3 0.3
Table 6: Hyperparameters specific to the embedding module, which generates the Wasserstein task embeddings online.
Hyperparameter Image Sequence Learning MiniGrid MiniHack
Reference size ($M$) 50 50 50
Num. samples ($N$) 128 128 128
Frequency (iterations) 1 1 1